OpenStudy (anonymous):

The elevation of the basement floor in a building is -14 ft. The elevation of the roof is 36 feet. What is the distance from the basement floor to the roof? A) -50 ft B) -24 ft C) 24 ft D) 50 ft

OpenStudy (anonymous):

Well basically all you do here is add 14 and 36 to get your answer. You finish :)

OpenStudy (wmj259):

y(final)-y(initial)=y(total) You want to solve for y(total). y(final) is the roof elevation. an y(initial) is your basement elevation.

OpenStudy (anonymous):

negative or positive? :/

OpenStudy (wmj259):

If you walk up the stairs what do you think the displacement should be?

OpenStudy (anonymous):

positive?

OpenStudy (wmj259):

Correct. Do you understand why it should be positive?

OpenStudy (anonymous):

The equation for this is | -14 + 36 | Solve

OpenStudy (anonymous):

because elavation is going up?

OpenStudy (anonymous):

Correct

OpenStudy (wmj259):

@TheBallisticOne, That does not look correct because you first are basically adding a negative number and then after taking the difference you take the absolute value.

OpenStudy (anonymous):

can i ask you guys some more? :)

OpenStudy (wmj259):

Based on the formula you gave that would mean the elevation is 22 feet.

OpenStudy (anonymous):

Oh you're right @wmj259 |-14|+36?

OpenStudy (anonymous):

i put 50

OpenStudy (anonymous):

i put 50

OpenStudy (anonymous):

Which is correct.

OpenStudy (anonymous):

kk next question :)

OpenStudy (anonymous):

The price of a gallon of gasoline was $1.65. The price rose y dollars. Then the price dropped $0.15 and later rose again by $0.45. Now the price of gasoline is? Which expression could be used to find the price of the gasoline? A) 1.9 + y Eliminate B) 1.95 + y C) 2.0 + y D) 2.1 + y

OpenStudy (anonymous):

First solve (1.65 - 0.15) + 0.45 The answer is (Above Solution) + y

OpenStudy (anonymous):

kk ill tell u my answer

OpenStudy (anonymous):

1.95+y?

OpenStudy (wmj259):

1.65+y-0.15+0.45

OpenStudy (anonymous):

Correct good job :)

OpenStudy (anonymous):

thxx :)

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

Which property explains why these two expressions are equal? 7 + q + 4 = 4 + 7 + q A) associative B) commutative C) distributive D) inverse

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Solve from these definitions Associative - Able to move around parenthesis and still get the same answer. Commutative - able to move numbers and variables and still get the same answer. Distributive - able to distribute a factor to two different numbers using the same property for instance k(7-9)=(k7)-(k9)

OpenStudy (anonymous):

commutative?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

WOW I cant believe i didnt get that _ . _

OpenStudy (anonymous):

A football player caught the ball and made a gain of 8 yards on the first play. On the next play the team lost 12 yards. How many yards do they need to gain on the next play to get back to their original position? A) 4 B) 6 C) 8 D) 12

OpenStudy (anonymous):

Because the numbers commute (that's how i remember)

OpenStudy (anonymous):

oh yea... that could help

OpenStudy (anonymous):

last question sorry for all the trouble :/

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

Assume the original position is 0.

OpenStudy (anonymous):

First play is the equation 0+8=

OpenStudy (anonymous):

which is 8

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Next play is the equation 8-12=

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

So how much do you need to add or subtract to get back to 0

OpenStudy (anonymous):

That's your answer

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

thanks for ALL the help :)

OpenStudy (anonymous):

You're very welcome :D

OpenStudy (anonymous):

wish i couldve gave you a medal i fanned you tho :)

OpenStudy (anonymous):

It's ok

OpenStudy (anonymous):

NOW i have to do a free response _ . _

OpenStudy (anonymous):

lol